OpenStudy (anonymous):

(1 pt) Monroe is graphing the equation . He has placed a point on the origin, and uses the slope to find two more points. Which points does Monroe plot on his graph? Choose exactly two answers that are correct. @aum @ganeshie8

OpenStudy (aum):

Are there some numbers missing from this problem?

OpenStudy (anonymous):

oh y=-4/9x

OpenStudy (aum):

slope = -4/9 slope = rise / run = -4/9 you can choose rise = -4 and run = +9 or rise = +4 and run = -9.

OpenStudy (anonymous):

A. (−9, 4) B. (−4, 9) C. (4, −9) D. (9, −4) @aum

OpenStudy (aum):

rise is along the y-axis and run is along the x-axis. start from the origin: case 1) rise = -4 and run = +9 x = 9, y = -4. Point is (9, -4) case 2) rise = +4 and run = -9 x = -9, y = 4. Point is (-9, 4)

OpenStudy (anonymous):

I need 2 anwers

OpenStudy (anonymous):

@au